# L2RSI: Cross-view LiDAR-based Place Recognition for Large-scale Urban Scenes via Remote Sensing Imagery

### v3

@@ Abstract
We tackle the challenge of LiDAR-based place recognition, which traditionally depends on costly and time-consuming prior 3D maps. To overcome this, we first construct XA-L&RSI dataset, which encompasses approximately $110,000$ remote sensing submaps and $13,000$ LiDAR point cloud submaps captured in urban scenes, and propose a novel method, L2RSI, for cross-view LiDAR place recognition using high-resolution Remote Sensing Imagery. This approach enables large-scale localization capabilities at a reduced cost by leveraging readily available overhead images as map proxies. L2RSI addresses the dual challenges of cross-view and cross-modal place recognition by learning feature alignment between point cloud submaps and remote sensing submaps in the semantic domain. Additionally, we introduce a novel probability propagation method based on particle estimation to refine position predictions, effectively leveraging temporal and spatial information. This approach enables large-scale retrieval and cross-scene generalization without fine-tuning. Extensive experiments on XA-L&RSI demonstrate that, within a $100km^{2}$ retrieval range, L2RSI accurately localizes $83.27\%$ of point cloud submaps within a $30m$ radius for top-$1$ retrieved location. Our project page is publicly available at https://shizw695.github.io/L2RSI/.
@@ 1 Introduction
Specific to the second problem, we propose the Spatial-Temproal Particle Estimation (STPE) to aggregate spatio-temporal information of continuous queries, thus to eliminate the ambiguity between geometrically distant but semantically similar positions. Specifically, the retrieval results in each query are considered as particles. Instead of filtering out the incorrect particles, our idea is to aggregate these particles and estimate a more accurate probability density distribution for the current position. Distribution of the particles of a single query can be modeled as the mixture of multiple gaussian models (GMMs). Then, the aggregation spread over a desired time window through an estimated inter-frame correspondence of the point clouds, estimating the probability density of the current query and refining the retrieval results.
@@ 4.2 Semantic contrastive learning
4.3 Spatial-Temproal Particle Estimation
@@ 5.1 Benchmark dataset
Due to the lack of publicly available datasets that provide correspondences between large-scale outdoor LiDAR scans and remote sensing imagery, we construct the XA-L&RSI and Oxford-L&RSI datasets for training and evaluation.
XA-L&RSI. We purchase high-resolution remote sensing imagery covering Xiang’an District from AIRSAT Technology Group, with a resolution of $0.5m$. We extract submaps of $60m\!\times\!60m$ from the original remote sensing imagery at intervals of $20m$, resulting in a two-third overlap between adjacent submaps. And the submaps lacking road semantics in their $30m\!\times\!30m$ central region are filtered out.
@@
Oxford-L&RSI. Oxford Radar RobotCar is a famous urban scene dataset [5] for localization.
(a) XA-L&RSI
(b) Oxford-L&RSI
Figure 3: Dataset configuration. In XA-L&RSI 3(a), different colored rectangles highlight test sets from databases of varying sizes, all sharing the same query trajectory (green). Besides, the training trajectory (light blue) and discarded trajectory at the boundary (red) are annotated. In Oxford-L&RSI 3(b), the green trajectory is as the query for the test set, while the entire remote sensing imagery serves as the database.
The data is collected with a Velodyne HDL-32E LiDAR. The dataset includes multiple repeated traversals, each covering about $10km$, as shown in Fig. 3(b). Based on this, we supplement the database with the corresponding remote sensing imagery, covering an area of approximately $2.8km^{2}$. And we simulate the magnetometer by adding random noise within $\pm 15^{\circ}$ to the ground truth of orientation, as the official dataset does not include a magnetometer. The source and resolution of remote sensing imagery, as well as the data preprocessing process, are the same as those of the XA-L&RSI. The data of 11-14-02-26, 14-12-05-52 are used as the test set to evaluate the generalization ability of the proposed method without fine-tuning, see Table 2.
@@
Submap Num. XA-L&RSI Oxford-L&RSI
@@ 5.3.1 Evaluation on test sets
Table 3: Performance comparison on XA-L&RSI and Oxford-L&RSI. $Recall@1$ and $Recall@10$ ($\%$) are reported. The best result is rendered in bold.
@@
The proposed L2RSI is the first to perform cross-view LiDAR-based place recognition in large-scale urban scenes with high-resolution remote sensing imagery. Several state-of-the-art place recognition networks are capable of handling preprocessed cross-modal semantic images, such as cross-modal 3D place recognition (LIP-Loc [21]) and cross-view 2D place recognition (GeoDTR [35], Sample4Geo [36], and FRGeo [37]). Firstly, we benchmark them to evaluate the performance of our method in the test sets of XA-L&RSI with different size databases. The training settings on XA-L&RSI for all methods are consistent, except for the image size and descriptor dimensions, which follow the official configurations. Notably, we replace the LiDAR range image projection used in the original LIP-Loc with a BEV projection.The comparison is summarized in Table 3. L2RSI (w. STPE) significantly outperforms the baseline methods, reaches a $Recall@1$ of $88.93\%$ with a database of $4km^{2}$. As the retrieval range expands, L2RSI (w. STPE) demonstrates increased robustness, with the database area increased to $100km^{2}$ and only a $5.66\%$ degradation in $Recall@1$ and a $4.31\%$ degradation in $Recall@10$. In addition, we also introduced a comparison of two refinement strategies, including SuperGlobal reranking [48] and classical particle filtering (PF). The former focuses on refining individual queries and database features, leading to only marginal improvements in performance. For the latter, we apply its principles within the retrieval framework, initializing the number of particles to $120$. With the integration of sequence information, the performance of L2RSI (w. PF) shows considerable improvement. However, in challenging cross-view and cross-modal retrieval tasks, particles frequently decay, necessitating re-initialization and impeding the flow of sequence information. The proposed STPE substitutes filtering with aggregation, efficiently utilizing spatio-temporal information to substantially enhance the unstable performance of single query.
Table 4: Ablation study about effectiveness of components. The results include $Recall@1$ and $Recall@30$ ($\%$) on Test Set-S and 11-14-02-26. The best result is rendered in bold. Sem.: Semantic segmentation. STPE: Spatial-Temproal Particle Estimation. Orient.: Orientation.
@@
Secondly, we directly test the model on the Oxford-L&RSI without any fine-tuning. Given the stark differences in equipment models, road widths, architectural styles, etc., this challenge marks the highest level of generalization ability. In the lower part of Table 3, the baseline methods fail completely, while L2RSI (STPE) achieves $Recall@1$ over $40\%$ and $Recall@10$ above $60\%$ on both trajectories. Although the current results are not yet sufficient for real-world deployment, they demonstrate great potential for future applications. Please refer to Sec. A.3 for comparative experiments in more scenes.
@@ 5.3.2 Ablation study
Semantic information. In the first row of Table 4, we bypass semantic images, directly extracting submaps from raw remote sensing imagery and projecting 3D coordinates to generate point cloud BEV images. Compared to the full L2RSI (last row), $Recall@1$ on test set-S decreases by $38.50\%$, and on Oxford-L&RSI, $Recall@1$ dropped to $8.13\%$. This demonstrates that semantics serve as reliable information capable of bridging domain differences, playing a pivotal role in enhancing the generalization ability.
@@
Figure 5: Visualization of retrieval results on XA-L&RSI 5(a) and Oxford-L&RSI 5(b). Points closer to red indicate that more retrieval results are required to achieve correct retrieval at that location. (I) shows the results without STPE, and (II) shows the results with STPE.
@@ 5.3.3 Impact of Motion Model Failure
where we take $\epsilon_{x}=\epsilon_{y}$ for convenience. The average Recall@1Recall@1($\%$) of three repeated experiments on Test Set-G ($100km^{2}$) of XA-L&RSI are shown in Table 6. Please note that the disturbances in the table represent intentionally introduced additional noise, based on the noise inherently caused by the FastGICP [38] algorithm and the magnetometer used in the system. $\epsilon_{yaw}=\pm 5^{\circ}$ and $\epsilon_{xy}=\pm 1$m indicates that, in addition to the inherent noise, random yaw angle noise within a range of $\pm 5^{\circ}$ and random translation noise within a range of $\pm 1m$ are introduced.
@@ 6 Conclusion
We present L2RSI, the first method for cross-view LiDAR-based place recognition within large-scale (over $100km^{2}$) urban scenes using high-resolution remote sensing imagery. We unify LiDAR point cloud submaps and remote sensing submaps into a shared semantic space via a carefully-designed contrastive learning network. Furthermore, we propose the Spatial-Temproal Particle Estimation to aggregate the spatio-temporal constraints of sequential queries, thereby further refining the performance of global place recognition. Extensive experiments on the XA-L&RSI and Oxford-L&RSI datasets demonstrate the effectiveness and generalization ability of L2RSI. Future work will focus on reducing the reliance on additional magnetometers and further addressing the challenges of deploying this technology in real-world applications.
@@ A.3 Comparison on more datasets
Table 8: Summary statistics for KITTI360-L&RSI.
@@
KITTI-360 is an urban scene dataset consisting of $9$ sequences with a mileage of approximately $80km$, commonly used for 3D perception and localization. To enable a more comprehensive evaluation, we additionally associate each trajectory with corresponding remote sensing imagery. Following the same procedure described in Sec. 5.1, we expand it to KITTI360-L&RSI, with six sequences for training (02, 03, 04, 05, 06, 07), one sequence for validation (00), and two sequences for testing (09, 10). The statistical information of the dataset is summarized in Table 8.
Table 9: Performance comparison on KITTI360-L&RSI. $Recall@1$ and $Recall@10$ ($\%$) are reported. The best result is rendered in bold.
@@ A.4 Impact of different semantic encoders
Table 10: Ablation study about the selection of the semantic encoder. $Recall@1$ and $Recall@10$ ($\%$) on XA-L&RSI and Oxford-L&RSI. $\dagger$} indicates that the input image size is $518\times 518$. The best result is rendered in bold.
@@
We consider several popular backbones in the selection of the semantic encoder, including VGG [51], ResNet [52], ConvNeXt [53], DINOv2 [54] and MAE [47]. Table 10 shows the results evaluated on the test sets of XA-L&RSI and Oxford-L&RSI. Specific pre-trained model versions are selected based on their GPU memory requirements. MAE-B performs the best, especially in terms of generalization ability. We attribute this to mask modeling during training, which aids the model in interpreting the sparse semantic images generated from point cloud projections. DINOv2-S from the official source uses a default image size of $518\times 518$. Images that are resized are marked with ($\dagger$}). Furthermore, to accommodate single-GPU training, the batch size is set to $8$, which hinder the model generalize to Oxford-L&RSI.
@@ A.5 Compatibility with existing pipelines
Table 11: Ablation study about the compatibility of STPE with other existing networks on XA-L&RSI and Oxford-L&RSI. $Recall@1$ and $Recall@10$ ($\%$) are reported. The best result is rendered in bold.
@@
The proposed STPE is easily compatible with existing place recognition pipelines. In Table 11, we refine the retrieval results of the baseline methods in using STPE. The results clearly indicate that all the methods achieve an improvement in recall on multiple test sets. However, when the vast majority of queries in the sequence fail completely, STPE is provided with insufficient useful information, as evidenced by the poor performance of Lip-Loc, GeoDTR and FRGeo on Oxford-L&RSI. Undoubtedly, when the relative position between LiDAR scans is available, the proposed STPE demonstrates superior effectiveness and compatibility with existing retrieval-based place recognition methods.

### v4

@@ Abstract
We tackle the challenge of LiDAR-based place recognition, which traditionally depends on costly and time-consuming prior 3D maps. To overcome this, we first construct LiRSI-XA dataset, which encompasses approximately $110,000$ remote sensing submaps and $13,000$ LiDAR point cloud submaps captured in urban scenes, and propose a novel method, L2RSI, for cross-view LiDAR place recognition using high-resolution Remote Sensing Imagery. This approach enables large-scale localization capabilities at a reduced cost by leveraging readily available overhead images as map proxies. L2RSI addresses the dual challenges of cross-view and cross-modal place recognition by learning feature alignment between point cloud submaps and remote sensing submaps in the semantic domain. Additionally, we introduce a novel probability propagation method based on particle estimation to refine position predictions, effectively leveraging temporal and spatial information. This approach enables large-scale retrieval and cross-scene generalization without fine-tuning. Extensive experiments on LiRSI-XA demonstrate that, within a $100km^{2}$ retrieval range, L2RSI accurately localizes $83.27\%$ of point cloud submaps within a $30m$ radius for top-$1$ retrieved location. Our project page is publicly available at https://shizw695.github.io/L2RSI/.
@@ 1 Introduction
Specific to the second problem, we propose the Spatial-Temporal Particle Estimation (STPE) to aggregate spatio-temporal information of continuous queries, thus to eliminate the ambiguity between geometrically distant but semantically similar positions. Specifically, the retrieval results in each query are considered as particles. Instead of filtering out the incorrect particles, our idea is to aggregate these particles and estimate a more accurate probability density distribution for the current position. Distribution of the particles of a single query can be modeled as the mixture of multiple gaussian models (GMMs). Then, the aggregation spread over a desired time window through an estimated inter-frame correspondence of the point clouds, estimating the probability density of the current query and refining the retrieval results.
@@ 4.2 Semantic contrastive learning
4.3 Spatial-Temporal Particle Estimation
@@ 5.1 Benchmark dataset
Due to the lack of publicly available datasets that provide correspondences between large-scale outdoor LiDAR scans and remote sensing imagery, we construct the LiRSI-XA and LiRSI-Oxford datasets for training and evaluation.
LiRSI-XA. We purchase high-resolution remote sensing imagery covering Xiang’an District from AIRSAT Technology Group, with a resolution of $0.5m$. We extract submaps of $60m\!\times\!60m$ from the original remote sensing imagery at intervals of $20m$, resulting in a two-third overlap between adjacent submaps. And the submaps lacking road semantics in their $30m\!\times\!30m$ central region are filtered out.
@@
LiRSI-Oxford. Oxford Radar RobotCar is a famous urban scene dataset [5] for localization.
(a) LiRSI-XA
(b) LiRSI-Oxford
Figure 3: Dataset configuration. In LiRSI-XA 3(a), different colored rectangles highlight test sets from databases of varying sizes, all sharing the same query trajectory (green). Besides, the training trajectory (light blue) and discarded trajectory at the boundary (red) are annotated. In LiRSI-Oxford 3(b), the green trajectory is as the query for the test set, while the entire remote sensing imagery serves as the database.
The data is collected with a Velodyne HDL-32E LiDAR. The dataset includes multiple repeated traversals, each covering about $10km$, as shown in Fig. 3(b). Based on this, we supplement the database with the corresponding remote sensing imagery, covering an area of approximately $2.8km^{2}$. And we simulate the magnetometer by adding random noise within $\pm 15^{\circ}$ to the ground truth of orientation, as the official dataset does not include a magnetometer. The source and resolution of remote sensing imagery, as well as the data preprocessing process, are the same as those of the LiRSI-XA. The data of 11-14-02-26, 14-12-05-52 are used as the test set to evaluate the generalization ability of the proposed method without fine-tuning, see Table 2.
@@
Submap Num. LiRSI-XA LiRSI-Oxford
@@ 5.3.1 Evaluation on test sets
Table 3: Performance comparison on LiRSI-XA and LiRSI-Oxford. $Recall@1$ and $Recall@10$ ($\%$) are reported. The best result is rendered in bold.
@@
The proposed L2RSI is the first to perform cross-view LiDAR-based place recognition in large-scale urban scenes with high-resolution remote sensing imagery. Several state-of-the-art place recognition networks are capable of handling preprocessed cross-modal semantic images, such as cross-modal 3D place recognition (LIP-Loc [21]) and cross-view 2D place recognition (GeoDTR [35], Sample4Geo [36], and FRGeo [37]). Firstly, we benchmark them to evaluate the performance of our method in the test sets of LiRSI-XA with different size databases. The training settings on LiRSI-XA for all methods are consistent, except for the image size and descriptor dimensions, which follow the official configurations. Notably, we replace the LiDAR range image projection used in the original LIP-Loc with a BEV projection.The comparison is summarized in Table 3. L2RSI (w. STPE) significantly outperforms the baseline methods, reaches a $Recall@1$ of $88.93\%$ with a database of $4km^{2}$. As the retrieval range expands, L2RSI (w. STPE) demonstrates increased robustness, with the database area increased to $100km^{2}$ and only a $5.66\%$ degradation in $Recall@1$ and a $4.31\%$ degradation in $Recall@10$. In addition, we also introduced a comparison of two refinement strategies, including SuperGlobal reranking [48] and classical particle filtering (PF). The former focuses on refining individual queries and database features, leading to only marginal improvements in performance. For the latter, we apply its principles within the retrieval framework, initializing the number of particles to $120$. With the integration of sequence information, the performance of L2RSI (w. PF) shows considerable improvement. However, in challenging cross-view and cross-modal retrieval tasks, particles frequently decay, necessitating re-initialization and impeding the flow of sequence information. The proposed STPE substitutes filtering with aggregation, efficiently utilizing spatio-temporal information to substantially enhance the unstable performance of single query.
Table 4: Ablation study about effectiveness of components. The results include $Recall@1$ and $Recall@30$ ($\%$) on Test Set-S and 11-14-02-26. The best result is rendered in bold. Sem.: Semantic segmentation. STPE: Spatial-Temporal Particle Estimation. Orient.: Orientation.
@@
Secondly, we directly test the model on the LiRSI-Oxford without any fine-tuning. Given the stark differences in equipment models, road widths, architectural styles, etc., this challenge marks the highest level of generalization ability. In the lower part of Table 3, the baseline methods fail completely, while L2RSI (STPE) achieves $Recall@1$ over $40\%$ and $Recall@10$ above $60\%$ on both trajectories. Although the current results are not yet sufficient for real-world deployment, they demonstrate great potential for future applications. Please refer to Sec. A.3 for comparative experiments in more scenes.
@@ 5.3.2 Ablation study
Semantic information. In the first row of Table 4, we bypass semantic images, directly extracting submaps from raw remote sensing imagery and projecting 3D coordinates to generate point cloud BEV images. Compared to the full L2RSI (last row), $Recall@1$ on test set-S decreases by $38.50\%$, and on LiRSI-Oxford, $Recall@1$ dropped to $8.13\%$. This demonstrates that semantics serve as reliable information capable of bridging domain differences, playing a pivotal role in enhancing the generalization ability.
@@
Figure 5: Visualization of retrieval results on LiRSI-XA 5(a) and LiRSI-Oxford 5(b). Points closer to red indicate that more retrieval results are required to achieve correct retrieval at that location. (I) shows the results without STPE, and (II) shows the results with STPE.
@@ 5.3.3 Impact of Motion Model Failure
where we take $\epsilon_{x}=\epsilon_{y}$ for convenience. The average Recall@1Recall@1($\%$) of three repeated experiments on Test Set-G ($100km^{2}$) of LiRSI-XA are shown in Table 6. Please note that the disturbances in the table represent intentionally introduced additional noise, based on the noise inherently caused by the FastGICP [38] algorithm and the magnetometer used in the system. $\epsilon_{yaw}=\pm 5^{\circ}$ and $\epsilon_{xy}=\pm 1$m indicates that, in addition to the inherent noise, random yaw angle noise within a range of $\pm 5^{\circ}$ and random translation noise within a range of $\pm 1m$ are introduced.
@@ 6 Conclusion
We present L2RSI, the first method for cross-view LiDAR-based place recognition within large-scale (over $100km^{2}$) urban scenes using high-resolution remote sensing imagery. We unify LiDAR point cloud submaps and remote sensing submaps into a shared semantic space via a carefully-designed contrastive learning network. Furthermore, we propose the Spatial-Temporal Particle Estimation to aggregate the spatio-temporal constraints of sequential queries, thereby further refining the performance of global place recognition. Extensive experiments on the LiRSI-XA and LiRSI-Oxford datasets demonstrate the effectiveness and generalization ability of L2RSI. Future work will focus on reducing the reliance on additional magnetometers and further addressing the challenges of deploying this technology in real-world applications.
@@ A.3 Comparison on more datasets
Table 8: Summary statistics for LiRSI-Kitti360.
@@
KITTI-360 [7] is an urban scene dataset consisting of $9$ sequences with a mileage of approximately $80km$, commonly used for 3D perception and localization. To enable a more comprehensive evaluation, we additionally associate each trajectory with corresponding remote sensing imagery. Following the same procedure described in Sec. 5.1, we expand it to LiRSI-Kitti360, with six sequences for training (02, 03, 04, 05, 06, 07), one sequence for validation (00), and two sequences for testing (09, 10). The statistical information of the dataset is summarized in Table 8.
Table 9: Performance comparison on LiRSI-Kitti360. $Recall@1$ and $Recall@10$ ($\%$) are reported. The best result is rendered in bold.
@@ A.4 Impact of different semantic encoders
Table 10: Ablation study about the selection of the semantic encoder. $Recall@1$ and $Recall@10$ ($\%$) on LiRSI-XA and LiRSI-Oxford. $\dagger$} indicates that the input image size is $518\times 518$. The best result is rendered in bold.
@@
We consider several popular backbones in the selection of the semantic encoder, including VGG [51], ResNet [52], ConvNeXt [53], DINOv2 [54] and MAE [47]. Table 10 shows the results evaluated on the test sets of LiRSI-XA and LiRSI-Oxford. Specific pre-trained model versions are selected based on their GPU memory requirements. MAE-B performs the best, especially in terms of generalization ability. We attribute this to mask modeling during training, which aids the model in interpreting the sparse semantic images generated from point cloud projections. DINOv2-S from the official source uses a default image size of $518\times 518$. Images that are resized are marked with ($\dagger$}). Furthermore, to accommodate single-GPU training, the batch size is set to $8$, which hinder the model generalize to LiRSI-Oxford.
@@ A.5 Compatibility with existing pipelines
Table 11: Ablation study about the compatibility of STPE with other existing networks on LiRSI-XA and LiRSI-Oxford. $Recall@1$ and $Recall@10$ ($\%$) are reported. The best result is rendered in bold.
@@
The proposed STPE is easily compatible with existing place recognition pipelines. In Table 11, we refine the retrieval results of the baseline methods in using STPE. The results clearly indicate that all the methods achieve an improvement in recall on multiple test sets. However, when the vast majority of queries in the sequence fail completely, STPE is provided with insufficient useful information, as evidenced by the poor performance of Lip-Loc, GeoDTR and FRGeo on LiRSI-Oxford. Undoubtedly, when the relative position between LiDAR scans is available, the proposed STPE demonstrates superior effectiveness and compatibility with existing retrieval-based place recognition methods.